ElCUHnningham:

Using a stopwatch, Tyrone determines it takes him 58.2 minutes to travel 30 miles to work. The stopwatch measures to hundredths of a minute. Going by the accuracy of the stopwatch, which is the most accurate determination for the number of feet per second Tyrone traveled on his way to work?

Zeke:

Are there choices??

Mercury:

we need to convert 30 miles/58.2 minutes to feet/second 1 mile = 5280 feet 1 minute = 60 seconds applying dimensional analysis \[\frac{ 30~miles }{ 52.8~minutes }\frac{ 1~minute }{ 60~seconds }\frac{ 5280~feet }{ 1~mile }\] *** important *** the question is asking to go by the significance of the stopwatch, so since 58.2 minutes has 3 sig figs, round your dimensional analysis to three digits

QuestionCoveBot:

The choices are shown below: 19.8 feet per second 19.84 feet per second 45.3 feet per second 45.36 feet per second